# Wasserstein Convergence of Score-based Generative Models under Semiconvexity and Discontinuous Gradients

### v1

@@ Abstract
Score-based Generative Models (SGMs) approximate a data distribution by perturbing it with Gaussian noise and subsequently denoising it via a learned reverse diffusion process. These models excel at modeling complex data distributions and generating diverse samples, achieving state-of-the-art performance across domains such as computer vision, audio generation, reinforcement learning, and computational biology. Despite their empirical success, existing Wasserstein-2 convergence analysis typically assume strong regularity conditions—such as smoothness or strict log-concavity of the data distribution—that are rarely satisfied in practice.
In this work, we establish the first non-asymptotic Wasserstein-2 convergence guarantees for SGMs targeting semiconvex distributions with potentially discontinuous gradients. Our upper bounds are explicit and sharp in key parameters, achieving optimal dependence of $O(\sqrt{d})$ on the data dimension $d$ and convergence rate of order one. The framework accommodates a wide class of practically relevant distributions, including symmetric modified half-normal distributions, Gaussian mixtures, double-well potentials, and elastic net potentials. By leveraging semiconvexity without requiring smoothness assumptions on the potential such as differentiability, our results substantially broaden the theoretical foundations of SGMs, bridging the gap between empirical success and rigorous guarantees in non-smooth, complex data regimes.
@@ 1 Introduction
Score-based Generative Models (SGMs), also known as diffusion-based generative models [55, 57, 54, 27], have rapidly emerged over the past few years as a popular approach in modern generative modelling due to their remarkable capabilities in generating complex data, surpassing previous state-of-the-art models, such as Generative Adversarial Networks (GANs) [22] and Variational AutoEncoders (VAEs) [34]. These models are now widely adopted in computer vision and audio generation tasks [35, 11, 47, 2, 32, 3, 53, 51, 72], text generation [41, 70, 45], sequential data modeling [1, 61, 62], reinforcement learning and control [49, 13, 24, 52, 73, 18], as well as life-science [14, 30, 64, 56, 65]. We refer the reader to the survey papers [69, 10] for a more comprehensive exposition of their applications.
@@
The primary goal of SGMs is to generate synthetic data that closely match a target data distribution $\pi_{\mathsf{D}}$, given a sample set. In particular, these models generate approximate data samples from high-dimensional data distributions by combining two diffusion processes, a forward and a backward process in time. The forward process is used to iteratively and smoothly transform samples from the unknown data distribution into (Gaussian) noise, while the associated backward process reverses the noising procedure to generate new samples from the starting unknown data distribution. A key role in these models is played by the score function, i.e. the gradient of the log-density of the solution of the forward process, which appears in the drift of the stochastic differential equation (SDE) associated with the backward process. Since this quantity depends on the unknown data distribution, an estimator of the score must be constructed during the noising step using score-matching techniques [28, 63].
@@
The widespread applicability and success of SGMs have been accompanied by a growing interest in the theoretical understandings of these models, particularly in the convergence analysis under different metrics such as Total Variation (TV) distance, Kullback Leibler (KL) divergence, Wasserstein distance, e.g., [5, 17, 6, 37, 67, 36, 44, 48, 38, 9, 12, 40, 50, 15, 4, 58, 60, 46, 21, 71]. In this work, we provide a non-asymptotic convergence analysis in Wasserstein distance of order two, as this metric is often considered more practical and informative for estimation tasks (see e.g., Equation (5)), and is closely connected to the popular Fréchet Inception Distance (FID) used to assess the quality of images in generative modeling (see, e.g., Section 4). A significant limitation of prior analysis in Wasserstein-2, e.g., [20, 7, 58, 60, 71], is their reliance on strong regularity conditions–such as smoothness or strict log-concavity – of the data distribution and its potential. These assumptions facilitate mathematical tractability but limit the applicability of theoretical results to more general settings, especially when the data distribution is only semiconvex and the potential’s gradient may be discontinuous. The only exception outside the strict log-concavity regime is the recent contribution in [21], where the authors assumes that the data distribution is weakly convex. However, their analysis still requires the potential to be twice continuously differentiable (see, e.g., [21, Proofs of Propositions B.1 and B.2]), and the stepsize of their generative algorithm must be bounded by a quantity inversely proportional to the one-sided Lipschitz constant of the potential (see [21, Equation (30)]). Still, such conditions on $\pi_{\mathsf{D}}$ in existing Wasserstein-2 convergence analysis do not fully reflect the complexity of real-world data, which often exhibit non-smooth or non-log-concave distributions. Therefore, the aim of this work is to address the following fundamental question:
@@
In addition, our estimates are explicit and exhibit the best known optimal dependencies in terms of data dimension, i.e., $O(\sqrt{d})$ in Theorem 3.12, and rate of convergence, i.e., $O(\gamma)$ in Theorem 3.14. In contrast to prior works under the same metric [21, 20, 58, 60], our estimates in Theorems 3.12 and Theorem 3.14 are derived without imposing any restrictions on the stepsize of the generative algorithm, making them more suitable to practical implementation. By circumventing the need for strict regularity conditions on the score function and allowing discontinuities in the gradients of the potentials, our work significantly expands the theoretical foundation of SGMs. This advancement not only bridges the gap between empirical success and theoretical guarantees but also opens new avenues for the application of diffusion models to data distributions with non-smooth potentials.
One source of error in the construction of the generative algorithm arises from replacing the initial condition of the backward process with the invariant measure of the forward process. To ensure this error remains small, the drift terms of both SDEs must satisfy, for instance, a monotonicity property with a time-dependent bound that meets an appropriate integrability condition (see, e.g., (19) and (23) below). To address this, we identify a time horizon for the generative algorithm that ensures the paths of the two backward processes become contractive. Notably, the integrability condition on the monotonicity bound depends only on the known constants in Assumption 2, making it significantly easier to verify in practice compared to the analogous condition in [21, Appendix C], which relies on weak convexity constants that are often difficult to estimate.
@@
1.1 Notation
[…]
$W_{2}(\mu,\nu)=\left(\inf_{\zeta\in\mathscr{C}(\mu,\nu)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|x-y|^{2}\ \text{d}\zeta(x,y)\right)^{\frac{1}{2}}.$
For any $x\in A\subseteq\mathbb{R}^{d}$ and any function $U:A\rightarrow\mathbb{R}$, the subdifferential $\partial U(x)$ of $U$ at $x$ is defined as
$\partial U(x)=\left\{p\in\mathbb{R}^{d}:\ \liminf_{y\rightarrow x}\frac{U(y)-U(x)-\langle p,y-x\rangle}{|y-x|}\geq 0\right\}.$ (1)
At the points where $U$ is differentiable, it holds that $\partial U(x)=\{\nabla U(x)\}$. For the sake of presentation, an element of $\partial U(x)$ is denoted by $h(x)$ for any $x\in\mathbb{R}^{d}$.
@@ 2 Technical Background for OU-based SGMs
In this section, we provide a brief summary behind the construction of score-based generative models based on diffusion introduced in [57]. The fundamental concept behind SGMs centers around the use of an ergodic (forward) diffusion process to diffuse the unknown data distribution $\pi_{\mathsf{D}}\in\mathscr{P}(\mathbb{R}^{d})$ to a known prior distribution and then learn a backward process to transform the prior to the target distribution $\pi_{\mathsf{D}}$ by estimating the score function of the forward process. In our analysis, we focus on the forward process $(X_{t})_{t\in[0,T]}$ given by an Ornstein-Uhlenbeck (OU) process as chosen in the original paper [57]
@@
where $(B_{t})_{t\in[0,T]}$ is an $d$-dimensional Brownian motion and we assume that $\mathbb{E}[|X_{0}|^{2}]<\infty$. Under mild assumptions on the target data distribution $\pi_{\mathsf{D}}$ [25, 8], the backward process $(Y_{t})_{t\in[0,T]}=(X_{T-t})_{t\in[0,T]}$ is given by
$\text{d}Y_{t}=(Y_{t}+2\nabla\log p_{T-t}(Y_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}\sim\mathscr{L}(X_{T}),$ (3)
where $\{p_{t}\}_{t\in[0,T]}$ is the family of densities of $\{\mathscr{L}(X_{t})\}_{t\in(0,T]}$ with respect to the Lebesgue measure, $\bar{B}_{t}$ is an another Brownian motion independent of $B_{t}$ in (2) defined on $(\Omega,\mathscr{F},\mathbb{P})$. However, the initial distribution in practice is the invariant measure of the forward process, which, in this case, is a standard $d$-dimensional Gaussian distribution. Therefore, the backward process (3) becomes
$\text{d}\widetilde{Y}_{t}=(\widetilde{Y}_{t}+2\ \nabla\log p_{T-t}(\widetilde{Y}_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad\widetilde{Y}_{0}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}).$ (4)
Since $\pi_{\mathsf{D}}$ is unknown, the score function $\nabla\log p_{t}$ in (3) cannot be computed exactly. To address this issue, an estimator $s(\cdot,\theta^{*},\cdot)$ is learned based on a family of functions $s:[0,T]\times\mathbb{R}^{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ parametrized in $\theta$, aiming at approximating the score of the ergodic (forward) diffusion process over a fixed time window $[0,T]$. In practice, the functions $s$ are neural networks and in particular cases, e.g. the motivating example in [7, Section 3.1], the functions $s$ can be wisely designed. The optimal value $\theta^{*}$ of the parameter $\theta$ is determined by optimizing the following score-matching objective
@@
An explicit expression of the stochastic gradient of (5) derived via denoising score matching [63] is provided in [7, Equation (8), Section 2]. Following [7, Section 2], we introduce an auxiliary process $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ containing the approximating function $s$ depending on the (random) estimator of $\theta^{*}$ denoted by $\hat{\theta}$, for $t\in[0,T]$,
$\text{d}Y_{t}^{\text{aux}}=(Y_{t}^{\text{aux}}+2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}^{\text{aux}}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}).$ (6)
The process (6) connects the backward process (4) and the numerical scheme (8), which facilitates the analysis of the convergence of the diffusion model (see the upper bounds in Appendix C involving $Y_{t}^{\text{aux}}$ for more details). Let the step size $\gamma_{k}=\gamma\in(0,1)$ for each $k=0,\dots,K$, where $K\in\mathbb{N}$. The discrete process $(Y_{k}^{\text{EM}})_{k\in\{0,\dots,K+1\}}$ of the Euler–Maruyama approximation of (6) is given, for any $k\in\{0,\dots,K\}$, as follows
$Y_{k+1}^{\text{EM}}=Y_{k}^{\text{EM}}+\gamma(Y_{k}^{\text{EM}}+2\ s(T-t_{k},\hat{\theta},Y_{k}^{\text{EM}}))+\sqrt{2\gamma}\ \bar{Z}_{k+1},\quad Y_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}),$ (7)
where $\{\bar{Z}_{k}\}_{k\in\{0,\dots,K+1\}}$ is a sequence of independent $d$-dimensional Gaussian random variables with zero mean and identity covariance matrix. The continuous-time interpolation of (7), for $t\in[0,T]$, is given by
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=(\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}+2\ s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}),$ (8)
where $\mathscr{L}(\widehat{Y}_{k}^{\text{EM}})=\mathscr{L}(Y^{\text{EM}}_{k})$ at grid points for each $k\in\{0,\dots,K+1\}$.

### v2

@@ Abstract
Score-based Generative Models (SGMs) approximate a data distribution by perturbing it with Gaussian noise and subsequently denoising it via a learned reverse diffusion process. These models excel at modeling complex data distributions and generating diverse samples, achieving state-of-the-art performance across domains such as computer vision, audio generation, reinforcement learning, and computational biology. Despite their empirical success, existing Wasserstein-2 convergence analysis typically assume strong regularity conditions–such as smoothness or strict log-concavity of the data distribution–that are rarely satisfied in practice. In this work, we establish the first non-asymptotic Wasserstein-2 convergence guarantees for SGMs targeting semiconvex distributions with potentially discontinuous gradients. Our upper bounds are explicit and sharp in key parameters, achieving optimal dependence of $O(\sqrt{d})$ on the data dimension $d$ and convergence rate of order one. The framework accommodates a wide class of practically relevant distributions, including symmetric modified half-normal distributions, Gaussian mixtures, double-well potentials, and elastic net potentials. By leveraging semiconvexity without requiring smoothness assumptions on the potential such as differentiability, our results substantially broaden the theoretical foundations of SGMs, bridging the gap between empirical success and rigorous guarantees in non-smooth, complex data regimes.
@@ 1 Introduction
Score-based Generative Models (SGMs), also known as diffusion-based generative models (Song & Ermon, 2019; Song et al., 2021; Sohl-Dickstein et al., 2015; Ho et al., 2020), have rapidly emerged over the past few years as a popular approach in modern generative modelling due to their remarkable capabilities in generating complex data, surpassing previous state-of-the-art models, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Variational AutoEncoders (VAEs) (Kingma & Welling, 2014). These models are now widely adopted in computer vision and audio generation tasks (Kong et al., 2020; Chen et al., 2021; Mittal et al., 2021; Avrahami et al., 2021; Kim et al., 2021; Bansal et al., 2023; Saharia et al., 2022; Po et al., 2023; Zhang et al., 2023), text generation (Li et al., 2022; Yu et al., 2022; Lovelace et al., 2023), sequential data modeling (Alcaraz & Strodthoff, 2023; Tashiro et al., 2021; Tevet et al., 2023), reinforcement learning and control (Pearce et al., 2023; Chi et al., 2023; Hansen-Estruch et al., 2023; Reuss et al., 2023; Zhu et al., 2023; Ding & Jin, 2024), as well as life-science (Chung & Ye, 2021; Jing et al., 2022; Watson et al., 2023; Song et al., 2022; Weiss et al., 2023). We refer the reader to the survey papers Yang et al. (2023); Chen et al. (2024) for a more comprehensive exposition of their applications.
@@
The primary goal of SGMs is to generate synthetic data that closely match a target data distribution $\pi_{\mathsf{D}}$, given a sample set. In particular, these models generate approximate data samples from high-dimensional data distributions by combining two diffusion processes, a forward and a backward process in time. The forward process is used to iteratively and smoothly transform samples from the unknown data distribution into (Gaussian) noise, while the associated backward process reverses the noising procedure to generate new samples from the starting unknown data distribution. A key role in these models is played by the score function, i.e. the gradient of the log-density of the solution of the forward process, which appears in the drift of the stochastic differential equation (SDE) associated with the backward process. Since this quantity depends on the unknown data distribution, an estimator of the score must be constructed during the noising step using score-matching techniques (Hyvärinen, 2005; Vincent, 2011).
@@
The widespread applicability and success of SGMs have been accompanied by a growing interest in the theoretical understandings of these models, particularly in the convergence analysis under different metrics such as Total Variation (TV) distance, Kullback Leibler (KL) divergence, Wasserstein distance, e.g., Block et al. (2020); De Bortoli et al. (2021); Bortoli (2022); Lee et al. (2022); Yang & Wibisono (2022); Kwon et al. (2022); Liu et al. (2022); Oko et al. (2023); Lee et al. (2023); Chen et al. (2023a; b); Li et al. (2024); Pedrotti et al. (2024); Conforti et al. (2025); Benton et al. (2024); Strasman et al. (2025); Bruno et al. (2025); Tang & Zhao (2024); Mimikos-Stamatopoulos et al. (2024); Wang & Wang (2024); Silveri & Ocello (2025); Yu & Yu (2025). In this work, we provide a non-asymptotic convergence analysis in Wasserstein distance of order two, as this metric is often considered more practical and informative for estimation tasks (see e.g., equation 4), and is closely connected to the popular Fréchet Inception Distance (FID) used to assess the quality of images in generative modeling (see, e.g., Section 4). A significant limitation of prior analysis in Wasserstein-2, e.g., Strasman et al. (2025); Gao et al. (2025); Bruno et al. (2025); Tang & Zhao (2024); Wang & Wang (2024); Yu & Yu (2025), is their reliance on strong regularity conditions–such as smoothness or strict log-concavity – of the data distribution and its potential. These assumptions facilitate mathematical tractability but limit the applicability of theoretical results to more general settings, especially when the data distribution is only semiconvex and the potential’s gradient may be discontinuous. The only exception outside the strict log-concavity regime is the recent contribution in Silveri & Ocello (2025), where the authors assumes that the data distribution is weakly convex. However, their analysis still requires the potential to be twice continuously differentiable (see, e.g., Silveri & Ocello (2025, Proofs of Propositions B.1 and B.2)), and the stepsize of their generative algorithm must be bounded by a quantity inversely proportional to the one-sided Lipschitz constant of the potential (see Silveri & Ocello (2025, equation (30))). Still, such conditions on $\pi_{\mathsf{D}}$ in existing Wasserstein-2 convergence analysis do not fully reflect the complexity of real-world data, which often exhibit non-smooth or non-log-concave distributions. Therefore, the aim of this work is to address the following fundamental question:
@@
In addition, our estimates are explicit and exhibit the best known optimal dependencies in terms of data dimension, i.e., $O(\sqrt{d})$ in Theorem 19, and rate of convergence, i.e., $O(\gamma)$ in Theorem 21. In contrast to prior works under the same metric Silveri & Ocello (2025); Gao et al. (2025); Strasman et al. (2025); Tang & Zhao (2024), our estimates in Theorem 19 and Theorem 21 are derived without imposing any restrictions on the stepsize of the generative algorithm [^1] The results in Silveri & Ocello (2025); Gao et al. (2025); Strasman et al. (2025); Tang & Zhao (2024) require the stepsize to be controlled in terms of the Lipschitz constant or the strong convexity constant of the target data distribution; see, e.g., Table 2 below. The only exceptions are the results in Bruno et al. (2025, Remark 12 and Theorem 10), which, however, requires stronger assumptions on the data distribution than Assumption 2 below., making them more suitable for practical implementation. By circumventing the need for strict regularity conditions on the score function and allowing discontinuities in the gradients of the potentials, our work significantly expands the theoretical foundation of SGMs. This advancement not only bridges the gap between empirical success and theoretical guarantees but also opens new avenues for the application of diffusion models to data distributions with non-smooth potentials.
One source of error in the construction of the generative algorithm arises from replacing the initial condition of the backward process with the invariant measure of the forward process. To ensure this error remains small, the drift terms of both SDEs must satisfy, for instance, a monotonicity property with a time-dependent bound that meets an appropriate integrability condition (see, e.g., equation 19 and equation 23 below). To address this, we identify a time horizon for the generative algorithm that ensures the paths of the two backward processes become contractive. Notably, the integrability condition on the monotonicity bound depends only on the known constants in Assumption 2, making it significantly easier to verify in practice compared to the analogous condition in Silveri & Ocello (2025, Appendix C), which relies on weak convexity constants that are often difficult to estimate.
@@
Notation. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a fixed probability space. We denote by $\mathbb{E}[X]$ the expectation of a random variable $X$. For $1\leq p<\infty$, $L^{p}$ is used to denote the usual space of $p$-integrable real-valued random variables. The $L^{p}$-integrability of a random variable $X$ is defined as $\mathbb{E}[|X|^{p}]<\infty$. Fix an integer $d\geq 1$. For an $\mathbb{R}^{d}$-valued random variable $X$, its law on $\mathcal{B}(\mathbb{R}^{d})$, i.e. the Borel sigma-algebra of $\mathbb{R}^{d}$ is denoted by $\mathcal{L}(X)$. Let $T>0$ denote a time horizon. For a positive real number $b$, we denote its integer part by $\lfloor b\rfloor$. The Euclidean scalar product is denoted by $\langle\cdot,\cdot\rangle$, with $|\cdot|$ standing for the corresponding norm (where the dimension of the space may vary depending on the context). Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuously differentiable function. The gradient of $f$ is denote by $\nabla f$. For any integer $q\geq 1$, let $\mathcal{P}(\mathbb{R}^{q})$ be the set of probability measures on $\mathcal{B}(\mathbb{R}^{q})$. For $\mu$, $\nu\in\mathcal{P}(\mathbb{R}^{d})$, let $\mathcal{C}(\mu,\nu)$ denote the set of probability measures $\zeta$ on $\mathcal{B}(\mathbb{R}^{2d})$ such that its respective marginals are $\mu$ and $\nu$. For any $\mu$ and $\nu\in\mathcal{P}(\mathbb{R}^{d})$, the Wasserstein distance of order 2 is defined as
$W_{2}(\mu,\nu)=\left(\inf_{\zeta\in\mathcal{C}(\mu,\nu)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|x-y|^{2}\ \text{d}\zeta(x,y)\right)^{\frac{1}{2}}.$
Table 4 (Appendix E) lists the main symbols used throughout this work along with references to where they are defined.
@@ 2 Technical Background for OU-based SGMs
In this section, we briefly summarize the construction of score-based generative models (SGMs) via diffusion processes, as introduced by Song et al. (2021). The core idea behind SGMs is to employ an ergodic (forward) diffusion process that gradually transforms the unknown data distribution $\pi_{\mathsf{D}}\in\mathcal{P}(\mathbb{R}^{d})$ into a known prior distribution. A backward (in time) process is then learned to transform the prior back to the target distribution $\pi_{\mathsf{D}}$ by estimating the score function of the forward process. In our analysis, we consider the forward process $(X_{t})_{t\in[0,T]}$ to be an Ornstein-Uhlenbeck (OU) process, consistent with the choice in the original paper Song et al. (2021)
@@
where $(B_{t})_{t\in[0,T]}$ is an $d$-dimensional Brownian motion and we assume that $\mathbb{E}[|X_{0}|^{2}]<\infty$.
For target data distributions $\pi_{\mathsf{D}}$ that are absolutely continuous with respect to the Lebesgue measure, and whose densities are continuous and integrable, the backward process $(Y_{t})_{t\in[0,T]}=(X_{T-t})_{t\in[0,T]}$ is well defined  [^2] The regularity of the Ornstein–Uhlenbeck semigroup for all $t\in(0,T]$ (see Appendix A, and Conforti et al. (2025, Proof of Proposition 3.1)) ensures that the necessary and sufficient conditions for the reversibility of the diffusion process are satisfied; see, e.g., Millet et al. (1989, Theorem 2.2) or (Haussmann & Pardoux, 1986, Theorem 2.1). These conditions on $\pi_{\mathsf{D}}$ are included in Assumption 2. (Millet et al., 1989; Haussmann & Pardoux, 1986), and is given by
$\text{d}Y_{t}=(Y_{t}+2\nabla\log p_{T-t}(Y_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}\sim\mathcal{L}(X_{T}),$ (2)
where $\{p_{t}\}_{t\in[0,T]}$ is the family of densities of $\{\mathcal{L}(X_{t})\}_{t\in(0,T]}$ with respect to the Lebesgue measure, $\bar{B}_{t}$ is an another Brownian motion independent of $B_{t}$ in 1 defined on $(\Omega,\mathcal{F},\mathbb{P})$. In practice, however, the initial distribution is taken to be the invariant measure of the forward process, which corresponds to the standard Gaussian distribution. As a result, the backward process in 2 becomes
$\text{d}\widetilde{Y}_{t}=(\widetilde{Y}_{t}+2\ \nabla\log p_{T-t}(\widetilde{Y}_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad\widetilde{Y}_{0}\sim\pi_{\infty}=\mathcal{N}(0,I_{d}).$ (3)
Since the target distribution $\pi_{\mathsf{D}}$ is unknown, the score function $\nabla\log p_{t}$ in 2 cannot be computed exactly. To overcome this limitation, an estimator $s(\cdot,\theta^{*},\cdot)$ is learned based on a family of functions $s:[0,T]\times\mathbb{R}^{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ parametrized in $\theta$, aiming at approximating the score of the ergodic forward process 4 over a fixed time window $[0,T]$. In practice, $s$ are neural networks and in particular cases, e.g., the motivating example in Bruno et al. (2025, Section 3.1), the functions $s$ can be carefully designed. The optimal value $\theta^{*}$ of the parameter $\theta$ is determined by optimizing the following score-matching objective
@@
An explicit expression of the stochastic gradient of 4 derived via denoising score matching (Vincent, 2011) is provided in Bruno et al. (2025, equation (8), Section 2). Following Bruno et al. (2025, Section 2), we define an auxiliary process $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ that incorporates the approximating function $s$, which depends on the (random) estimator of $\theta^{*}$ denoted by $\hat{\theta}$. For $t\in[0,T]$, this process is given by
$\text{d}Y_{t}^{\text{aux}}=(Y_{t}^{\text{aux}}+2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}^{\text{aux}}\sim\pi_{\infty}=\mathcal{N}(0,I_{d}).$ (5)
The auxiliary process 5 serves as a bridge between the backward process 3 and the numerical scheme 7, and it facilitates the analysis of the convergence of the diffusion model (see the upper bounds involving $Y_{t}^{\text{aux}}$ in the proof of Theorem 19 in Appendix C for further details). We now introduce the numerical scheme. Let the step size $\gamma_{j}=\gamma\in(0,1)$ for each $j=0,\dots,J$, where $J\in\mathbb{N}$. The discrete process $(Y_{j}^{\text{EM}})_{j\in\{0,\dots,J+1\}}$ of the Euler–Maruyama approximation of 5 is given, for any $j\in\{0,\dots,J\}$, as follows
$Y_{j+1}^{\text{EM}}=Y_{j}^{\text{EM}}+\gamma(Y_{j}^{\text{EM}}+2\ s(T-t_{j},\hat{\theta},Y_{j}^{\text{EM}}))+\sqrt{2\gamma}\ \bar{Z}_{j+1},\quad Y_{0}^{\text{EM}}\sim\pi_{\infty}=\mathcal{N}(0,I_{d}),$ (6)
where $\{\bar{Z}_{j}\}_{j\in\{0,\dots,J+1\}}$ is a sequence of independent $d$-dimensional Gaussian random variables with zero mean and identity covariance matrix. The continuous-time interpolation of 6, for $t\in[0,T]$, is given by
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=(\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}+2\ s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathcal{N}(0,I_{d}),$ (7)
where $\mathcal{L}(\widehat{Y}_{j}^{\text{EM}})=\mathcal{L}(Y^{\text{EM}}_{j})$ at grid points for each $j\in\{0,\dots,J+1\}$.